OpenStudy (mellamokatie):

Simplify the expression (3x/2)^4 A. 81x^4/16 B. 6x^4 C. 12x^4/2 D. 81x^4/2

OpenStudy (mellamokatie):

@kittiwitti1 When you're done can you help me please? Thanks

OpenStudy (kittiwitti1):

This?\[\left(\frac{3x}{2}\right)^{4}\]

OpenStudy (mellamokatie):

Yes :D

OpenStudy (kittiwitti1):

Er... I am not sure. I have an idea but it may be wrong D:

OpenStudy (mellamokatie):

hmm yeah I am lost haha

OpenStudy (mellamokatie):

I have to go I'll be back my sister needs the computer D:

OpenStudy (kittiwitti1):

Alright

OpenStudy (kittiwitti1):

https://goo.gl/GTnFCr This might help

OpenStudy (naka354):

(3x/2)^4 = ((3x)^4)/(2^4) = (81x^4)/16

OpenStudy (mellamokatie):

Thank you!

OpenStudy (kittiwitti1):

-.-